# SpeechPrune: Context-aware Token Pruning for Speech Information Retrieval

### v1

@@ V Conclusion
In this work, we introduced the SIR task to target long-form speech comprehension, presented SPIRAL as a benchmark for evaluating such capabilities, and proposed SpeechPrune, a training-free token pruning method leveraging speech-text similarity and approximate attention. Experimental results showed that SpeechPrune not only reduces computational costs but can also enhance model performance, achieving nearly 29% and up to 47% network accuracy improvements over the original model and the random pruning model, respectively. While promising, further exploration is needed to improve robustness to diverse audio conditions, explore more nuanced token selection methods, and adapt pruning strategies to specific input characteristics or fine-tuned models.

### v2

@@ V Conclusion
In this work, we introduced the SIR task to target long-form speech comprehension, presented SPIRAL as a benchmark for evaluating such capabilities, and proposed SpeechPrune, a training-free token pruning method leveraging speech-text similarity and approximate attention. Experimental results showed that SpeechPrune not only reduces computational costs but can also enhance model performance, achieving network accuracy improvements of nearly $29\%$ and up to $47\%$ over the original model and the random pruning model, respectively. While promising, further exploration is needed to improve robustness under diverse audio conditions, explore additional token selection methods, and adapt pruning strategies to specific input characteristics or fine-tuned models.